A 40-pound weight is placed 6 feet from the fulcrum on a lever, and a 60-pound weight is placed 8 feet from the fulcrum on the other side.
Where should a 30-pound weight be placed to give equilibrium?
This is a question on a handout and I can't find the formula to work this problem.
Any help would be apprecitated.
Thanks so much.
: A 40-pound weight is placed 6 feet from the fulcrum on a lever, and a 60-pound weight is placed 8 feet from the fulcrum on the other side.
Where should a 30-pound weight be placed to give equilibrium?
This is a question on a handout and I can't find the formula to work this problem.
Any help would be apprecitated.
Thanks so much.

Answer by scott8148(2761)

the weight multiplied by the distance from the fulcrum is a quantity called the "moment"
__ at equilibrium, the moments on either side of the fulcrum are equal

40(6)+30(x)=60(8) __ 240+30x=480 __ 30x=240 __ x=8